a. One set of 10 pairs of scores, X and Y values, produces a correlation of r = 0.60. If SSY = 200, find the standard error of estimate for the regression line.
b. A second set of 10 pairs of X and Y values produces a correlation of r = 0.40. If SSY = 200, find the standard error of estimate for the regression line.
Answer #1

a) standard error of estimate =sqrt((1-r2)*SSY/(n-2))=sqrt((1-0.62)*200/8)=4

b)

standard error of estimate =sqrt((1-0.42)*200/8)=4.5826